# Optimizing VarLiNGAM for Scalable and Efficient Time Series Causal Discovery

### v1

@@ Abstract
Causal discovery is designed to identify causal relationships in data, a task that has become increasingly complex due to the computational demands of traditional methods such as VarLiNGAM, which combines Vector Autoregressive Model with Linear Non-Gaussian Acyclic Model for time series data.
This study is dedicated to optimising causal discovery specifically for time series data, which is common in practical applications. Time series causal discovery is particularly challenging due to the need to account for temporal dependencies and potential time lag effects. By designing a specialised dataset generator and reducing the computational complexity of the VarLiNGAM model from $O(m^{3}\cdot n)$ to $O(m^{3}+m^{2}\cdot n)$, this study significantly improves the feasibility of processing large datasets. The proposed methods have been validated on advanced computational platforms and tested across simulated, real-world, and large-scale datasets, showcasing enhanced efficiency and performance. The optimised algorithm achieved 7 to 13 times speedup compared with the original algorithm and around 4.5 times speedup compared with the GPU-accelerated version on large-scale datasets with feature sizes between 200 and 400.
Our methods aim to push the boundaries of current causal discovery capabilities, making them more robust, scalable, and applicable to real-world scenarios, thus facilitating breakthroughs in various fields such as healthcare and finance.
Keywords: Causal Discovery, Time series, VarLiNGAM, Dataset Generator, Precomputation, Computational Complexity, Entropy.
@@
Causal discovery can be applied to both non-temporal data and time series data. Non-temporal data are data that is collected at a specific moment and is characterised by the fact that the data does not change over time, or is analysed with the assumption that the data does not change over time. Time series is a form of data in which data points are arranged in chronological order, usually recorded at equal intervals. It contains information about changes in variables over time, and these changes may contain important causal information.
@@
    To test the general applicability of the optimised algorithm, we applied it to large datasets generated by the previously mentioned generators, real-world datasets with ground truth, and large-scale real datasets without ground truth. The results before and after optimisation were compared to verify the accuracy and efficiency of the algorithm. Further details are given in Chapter 5.
@@ 2.1 Time Series Data
Time series data, observed in chronological order, are generated across various domains such as healthcare and finance. In medical studies, time series data are used to examine the impact of new treatments or drug interventions over time. For instance, time series data are studied to determine how the introduction of a new medication affects patient outcomes like blood pressure or recovery rates [47]. In finance, time series data allow analysts to study the effects of monetary policies or market interventions on stock prices, inflation, or interest rates. For instance, time series data are analysed to evaluate how central bank interest rate changes influence stock market performance [2].
Over the years, different tasks such as classification, clustering, and prediction have been proposed to analyze these data. One common classification task in finance is predicting whether a stock price will increase or decrease based on historical data. For example, algorithms like Random Forest or Support Vector Machines (SVMs) are trained using features such as past stock prices, trading volumes, and economic indicators. A specific example of this is using time series data to classify whether the stock of a company will rise or fall the next day based on a variety of indicators from the previous days [6]. In healthcare, predictive models help forecast patient outcomes. For instance, time series data collected from vital signs (such as heart rate, blood pressure, and oxygen levels) of patients can be used to predict the likelihood of a patient being readmitted to the hospital after discharge. Deep learning models like Long Short-Term Memory (LSTM) networks have been utilised to anticipate the onset of heart failure based on patient history, improving early detection [12].
@@ 2.2 Causal Discovery for Non-Temporal Data
This section provides an overview of two basic methods for causal discovery in non-temporal data, focusing on the fundamental modelling ideas of the relevant methods.
@@ 2.3 Causal Discovery for Time Series Data
This section gives an overview of the fundamentals of observational data-based methods for causal inference for time series, focusing on the modelling ideas of the relevant methods as well as the differences and connections between methods of the same category and different branches. Table 2.1 provides a general classification of some popular methods.
@@ 2.3.1 Constraint-based methods
Neapolitan [39] proposed Bayesian network (BN), which provides a new effective theoretical model for uncertain knowledge representation and reasoning. Causal Bayesian Network (causal BN, CBN) is the basis of causal network structure learning algorithms, which further specifies based on BN that directed edges represent causal relationships, out-degree nodes represent cause variables, and in-degree nodes represent effect variables. The expansion from BN to CBN usually requires the fulfilment of the fidelity assumption and the causal Markov condition [54], among others.
@@ 2. Methods based on information theory
3. Methods based on structural causal modelling
The earliest approach based on structural causal modelling is the linear non-Gaussian acyclic model (LiNGAM) proposed by Shimizu et al. [51]. Shimizu et al. [52] also proposed the nonparametric DirectLiNGAM framework. DirectLiNGAM estimates the causal order by identifying the exogenous variables and controls the false positives of causal inference in the multivariate case. As the sample size increases, DirectLiNGAM guarantees asymptotic convergence to the correct solution in a fixed number of small steps.
The LiNGAM method relies on non-Gaussian assumptions about the distribution of noise, which poses a significant limitation to the application of the model. Hoyer et al. [27] proposed the additive noise model (ANM), which assumes that the noise acts in the form of additive terms between the dependent variables, it does not need to rely on non-Gaussian assumptions about the noise and is not limited to linear regression. In addition, ANM determines the causal direction by utilising methods such as subsequent independence test [37] or class entropy score [40], in addition to residual independence test. Zhang and Hyvarinen [59] proposed post non-linear causal model (PNLCM). PNLCM builds a composite nonlinear model based on LiNGAM by performing two nonlinear transformations of the cause variable and the noise term. Through vectorization and matrix computation, LiNGAM, ANM and PNLCM are easily extended to the case of multiple variables. In addition, time-lagged causal inference of time series can be realised by data panning.
@@ 4. Methods based on nonlinear state space models
In 2012, Sugihara et al. [55] proposed the convergent cross mapping (CCM) method, which infer causality through state-space reconstruction. The method based on nonlinear state space modelling has been applied in neuroscience as early as the beginning of the 20th century, and in recent years, it has been widely used in many fields such as ecology and meteorological science, with the most applications in large complex systems.
State space modelling is an effective form of describing and revealing the internal relationships and laws of motion of a system. The time series causal inference method based on the nonlinear state space model assumes that the interaction occurs in a potential dynamical system, and then the causal relationship is inferred based on Takens’ theorem and nonlinear state space reconstruction. Takens’ theorem can be used to reconstruct the dynamical information in a time series, and it proves that the mapping from an attractor to the reconstruction space is one-to-one when certain conditions are satisfied, and that the reconstruction of orbitals in the phase space can be achieved by finding a reasonable number of embedding dimensions and keeping their original differential structure unchanged.
@@ 2.4.2 Acceleration for function-based method
Deng et al. [14] proposed an approach that integrates deep learning techniques into causal discovery to handle high-dimensional data and complex relationships. This integration addresses traditional bottlenecks in causal learning, such as high-dimensional variables and combinatorial optimisation problems. Deep neural networks contribute to causal learning by addressing conventional challenges from three aspects: representation, discovery, and inference.
MATSUDA et al. [36] proposed a supercomputer-accelerated approach to enhance the performance of LiNGAM causal discovery. The method leverages SIMD (Single Instruction, Multiple Data) instructions and MPI (Message Passing Interface) parallelization to overcome the computational complexity bottleneck of LiNGAM, which is $O(d^{3})$. Initially, SIMD instructions are utilised to parallelize mathematical operations, significantly speeding up the computation of mutual information differences and regression analyses. Subsequently, MPI parallelization is employed to distribute the computation across multiple nodes of the Fugaku supercomputer. This method effectively leverages Fugaku’s computational power and memory, allowing the handling of over 10,000 variables. Evaluations on 96 nodes showed a 17,531-fold speed-up over the original Python implementation, completing 20,000-variable computations in 17.7 hours. This approach significantly enhances the accuracy-speed trade-off compared to CPU and GPU-based tools.
@@ 2.5 Time-Series Causal Discovery
The output of time series causal discovery algorithms are causal graphs [4]. These graphs could help researchers distinguish between causal and correlational relationships, support causal inference, and identify critical causal paths in complex systems. By visually representing causal relationships, it helps to understand and predict system behaviour, guide experimental design and improve decision-making, leading to more accurate analysis and more reliable predictions.
@@ 2.6 Summary
This chapter introduces the basic concepts and methods for causal discovery in both non-temporal and time-series data, as well as techniques for analysing and accelerating these methods. First, it introduces time-series data to establish a basic understanding of dataset format. The chapter then explores causal discovery in non-temporal series data, starting with constraint-based approaches, which use conditions such as conditional independence tests to infer causality, and then function-based approaches, which model causal structure through functional relationships. After that, the chapter focuses on the causal discovery for time series data, using a similar approach but taking into account the time dependence and time lags in such data. Then, it discusses acceleration techniques, presenting strategies for optimising constraint-based and function-based approaches to handle larger datasets and improve computational efficiency. Finally, the chapter summarises the differences between time-series causal discovery and non-temporal causal discovery, as well as the structure of three types of causal graphs, providing a foundation for subsequent discussions.
@@ 4.1 Speed Bottleneck of VarLiNGAM
VarLiNGAM [31] is a causal inference method for time series data with a process that combines the ideas of autoregressive modelling with LiNGAM. The workflow of VarLiNGAM is as follows:
@@ 4. Estimate the causal effect matrix $B_{\tau}$ by the following equation:
By using a performance analysis tool, it is possible to roughly analyse the time spent on each step throughout the model run. Among all these procedures, estimating the VAR coefficients and calculating the residual vectors only takes less than 1% of the total execution time, whereas the most time-consuming procedure is to perform a LiNGAM analysis to find the $B_{0}$ matrix, which takes over half of the total execution time when the dataset is large.
@@ 4.1.1 Time Complexity Analysis of LiNGAM
By using the performance analysis tool, it is possible to roughly analyse the time spent on each step throughout DirectLiNGAM model run. The main computational bottleneck of the DirectLiNGAM algorithm is _search_causal_order, which takes over 96% of the overall execution time.
@@ 4.2.1 Time Complexity After Precomputation
The idea of using pre-computed entropies in the _search_causal_order process is motivated by a common optimisation strategy in algorithm design called avoiding redundant calculations. This approach is often referred to as memoization or precomputation, and it is particularly useful in scenarios where the same computationally expensive operations are repeatedly performed.
@@ 4.3 Summary
This chapter focuses on optimising the VarLiNGAM algorithm by addressing its main speed bottlenecks and proposing acceleration strategies. First, it identifies that the most time-consuming process in VarLiNGAM is the LiNGAM analysis, specifically the step that searches for causal order. This step can take up a significant portion of the total execution time, especially for large datasets. To overcome this, the chapter introduces a precomputation strategy, which reduces redundant calculations by precomputing entropies and residual entropies before the causal order search. This significantly reduces the overall time complexity from $O(m^{3}\cdot n)$ to $O(m^{3}+m^{2}\cdot n)$, making the algorithm more efficient. Additionally, it also highlights the trade-off between time and space complexity, as precomputation increases memory usage. Nevertheless, these optimisations make VarLiNGAM still well-suited for large-scale datasets.
@@ 5.2.2 Validation: Fixed sample points
The graph 5.1(c) shows a sharp increase with larger feature sizes, similar to the trend seen in 5.1(a). This reinforces the observation that as feature size increases, the overall processing time becomes increasingly prohibitive without optimisation techniques like precomputation.
@@ 5.2.3 Validation: Fixed feature size
As seen in 5.2(e) and 5.2(f), the performance of the causal graph improves slightly as the number of sample points increases and stabilises after about 500 sample points have been considered. Precomputation appears to have a stabilising effect on the performance metrics. This indicates that the algorithm may be more robust after optimisation, maintaining its effectiveness across a broader range of data sizes. The consistency of the F1-score, Precision, Recall, and Accuracy after precomputation suggests that the capability of the algorithm is preserved despite the optimisation.
@@ 5.3.1 Data from Simulation: fMRI
The minor reduction in the F1-score suggests that the precomputation does not significantly affect the performance of the model. This small difference is within the standard deviation range, indicating that the predictive performance remains stable even after optimisation.
@@ 5.3.2 Real Dataset: IT Monitoring Datasets
For Web1, the F1-score decreases slightly after precomputation, indicating a minimal reduction in predictive performance. The decrease is very small, suggesting that the optimisation did not significantly affect the model’s ability to classify or predict. The total execution time decreases from 2.588 seconds to 2.411 seconds after precomputation, despite the additional precomputation step. The significant reduction in the search causal order time (from 0.374 to 0.011 seconds) contributes to the overall efficiency improvement. The decrease in execution time is around 6.8%.
@@
The results for these datasets show that precomputation can significantly optimise time complexity without compromising, and in some cases slightly improving the performance. This makes precomputation a valuable optimisation technique, especially for scenarios where large datasets are involved and quick execution is crucial.
@@ 6.1 Conclusion
This study addresses the challenges associated with causal discovery in time series data, particularly in the context of high-dimensional datasets. By optimising the VarLiNGAM model, we have significantly reduced its computational complexity, making it feasible to handle large-scale datasets that were previously unmanageable.
Our work began with a comprehensive review of causal inference methods for both non-temporal and time series data, where we analyzed the performance of constraint-based and function-based approaches on complex datasets. We chose to focus on the VarLiNGAM model and designed a specialised dataset generator to generate simulated large-scale data. After that, we used some practical tools to identify the key bottlenecks in its computational processes. Through analysis of the whole process, we proposed acceleration methods tailored to these challenges by using precomputation on the entropies of each feature and the residuals of each pair. Specifically, we investigated the use of precomputation to streamline the execution of time-intensive functions, including _search_causal_order and _diff_mutual_info. The time complexity analysis shows that precomputation can significantly reduce the time complexity of the algorithm, a conclusion that is strongly supported by experimental results on a variety of datasets including simulated datasets, real fMRI datasets, IT monitoring datasets, and the SP500 financial dataset.

### v2

@@ Abstract
Causal discovery identifies causal relationships in data, but the task is more complex for multivariate time series due to the computational demands of methods like VarLiNGAM, which combines a Vector Autoregressive Model with a Linear Non-Gaussian Acyclic Model. This study optimizes causal discovery specifically for time series data, which are common in practical applications. Time series causal discovery is particularly challenging because of temporal dependencies and potential time lag effects. By developing a specialized dataset generator and reducing the computational complexity of the VarLiNGAM model from $O(m^{3}\cdot n)$ to $O(m^{3}+m^{2}\cdot n)$, this study enhances the feasibility of processing large datasets. The proposed methods were validated on advanced computational platforms and tested on simulated, real-world, and large-scale datasets, demonstrating improved efficiency and performance. The optimized algorithm achieved 7 to 13 times speedup compared to the original and about 4.5 times speedup compared to the GPU-accelerated version on large-scale datasets with feature sizes from 200 to 400. Our methods extend current causal discovery capabilities, making them more robust, scalable, and applicable to real-world scenarios, facilitating advancements in fields like healthcare and finance.
@@
Causal discovery can be applied to both non-temporal data and time series data. Non-temporal data are data that is collected at a specific moment and is characterised by the fact that the data does not change over time, or is analyzed with the assumption that the data does not change over time. Time series is a form of data in which data points are arranged in chronological order, usually recorded at equal intervals. It contains information about changes in variables over time, and these changes may contain important causal information.
@@
    To test the general applicability of the optimised algorithm, we applied it to large datasets generated by the previously mentioned generators, real-world datasets with ground truth, and large-scale real datasets without ground truth. The results before and after optimization were compared to verify the accuracy and efficiency of the algorithm. Further details are given in Chapter 5.
@@ 2.1 Time Series Data
Time series data, observed in chronological order, are generated across various domains such as healthcare and finance. In medical studies, time series data are used to examine the impact of new treatments or drug interventions over time. For instance, time series data are studied to determine how the introduction of a new medication affects patient outcomes like blood pressure or recovery rates [47]. In finance, time series data allow analysts to study the effects of monetary policies or market interventions on stock prices, inflation, or interest rates. For instance, time series data are analyzed to evaluate how central bank interest rate changes influence stock market performance [2].
Over the years, different tasks such as classification, clustering, and prediction have been proposed to analyze these data. One common classification task in finance is predicting whether a stock price will increase or decrease based on historical data. For example, algorithms like Random Forest or Support Vector Machines (SVMs) are trained using features such as past stock prices, trading volumes, and economic indicators. A specific example of this is using time series data to classify whether the stock of a company will rise or fall the next day based on a variety of indicators from the previous days [6]. In healthcare, predictive models help forecast patient outcomes. For instance, time series data collected from vital signs (such as heart rate, blood pressure, and oxygen levels) of patients can be used to predict the likelihood of a patient being readmitted to the hospital after discharge. Deep learning models like Long Short-Term Memory (LSTM) networks have been utilized to anticipate the onset of heart failure based on patient history, improving early detection [12].
@@ 2.2 Causal Discovery for Non-Temporal Data
This section provides an overview of two basic methods for causal discovery in non-temporal data, focusing on the fundamental modeling ideas of the relevant methods.
@@ 2.3 Causal Discovery for Time Series Data
This section gives an overview of the fundamentals of observational data-based methods for causal inference for time series, focusing on the modeling ideas of the relevant methods as well as the differences and connections between methods of the same category and different branches. Table 2.1 provides a general classification of some popular methods.
@@ 2.3.1 Constraint-based methods
Neapolitan [39] proposed Bayesian network (BN), which provides a new effective theoretical model for uncertain knowledge representation and reasoning. Causal Bayesian Network (causal BN, CBN) is the basis of causal network structure learning algorithms, which further specifies based on BN that directed edges represent causal relationships, out-degree nodes represent cause variables, and in-degree nodes represent effect variables. The expansion from BN to CBN usually requires the fulfillment of the fidelity assumption and the causal Markov condition [54], among others.
@@ 2. Methods based on information theory
3. Methods based on structural causal modeling
The earliest approach based on structural causal modeling is the linear non-Gaussian acyclic model (LiNGAM) proposed by Shimizu et al. [51]. Shimizu et al. [52] also proposed the nonparametric DirectLiNGAM framework. DirectLiNGAM estimates the causal order by identifying the exogenous variables and controls the false positives of causal inference in the multivariate case. As the sample size increases, DirectLiNGAM guarantees asymptotic convergence to the correct solution in a fixed number of small steps.
The LiNGAM method relies on non-Gaussian assumptions about the distribution of noise, which poses a significant limitation to the application of the model. Hoyer et al. [27] proposed the additive noise model (ANM), which assumes that the noise acts in the form of additive terms between the dependent variables, it does not need to rely on non-Gaussian assumptions about the noise and is not limited to linear regression. In addition, ANM determines the causal direction by utilising methods such as subsequent independence test [37] or class entropy score [40], in addition to residual independence test. Zhang and Hyvarinen [59] proposed post non-linear causal model (PNLCM). PNLCM builds a composite nonlinear model based on LiNGAM by performing two nonlinear transformations of the cause variable and the noise term. Through vectorization and matrix computation, LiNGAM, ANM and PNLCM are easily extended to the case of multiple variables. In addition, time-lagged causal inference of time series can be realized by data panning.
@@ 4. Methods based on nonlinear state space models
In 2012, Sugihara et al. [55] proposed the convergent cross mapping (CCM) method, which infer causality through state-space reconstruction. The method based on nonlinear state space modeling has been applied in neuroscience as early as the beginning of the 20th century, and in recent years, it has been widely used in many fields such as ecology and meteorological science, with the most applications in large complex systems.
State space modeling is an effective form of describing and revealing the internal relationships and laws of motion of a system. The time series causal inference method based on the nonlinear state space model assumes that the interaction occurs in a potential dynamical system, and then the causal relationship is inferred based on Takens’ theorem and nonlinear state space reconstruction. Takens’ theorem can be used to reconstruct the dynamical information in a time series, and it proves that the mapping from an attractor to the reconstruction space is one-to-one when certain conditions are satisfied, and that the reconstruction of orbitals in the phase space can be achieved by finding a reasonable number of embedding dimensions and keeping their original differential structure unchanged.
@@ 2.4.2 Acceleration for function-based method
Deng et al. [14] proposed an approach that integrates deep learning techniques into causal discovery to handle high-dimensional data and complex relationships. This integration addresses traditional bottlenecks in causal learning, such as high-dimensional variables and combinatorial optimization problems. Deep neural networks contribute to causal learning by addressing conventional challenges from three aspects: representation, discovery, and inference.
MATSUDA et al. [36] proposed a supercomputer-accelerated approach to enhance the performance of LiNGAM causal discovery. The method leverages SIMD (Single Instruction, Multiple Data) instructions and MPI (Message Passing Interface) parallelization to overcome the computational complexity bottleneck of LiNGAM, which is $O(d^{3})$. Initially, SIMD instructions are utilized to parallelize mathematical operations, significantly speeding up the computation of mutual information differences and regression analyzes. Subsequently, MPI parallelization is employed to distribute the computation across multiple nodes of the Fugaku supercomputer. This method effectively leverages Fugaku’s computational power and memory, allowing the handling of over 10,000 variables. Evaluations on 96 nodes showed a 17,531-fold speed-up over the original Python implementation, completing 20,000-variable computations in 17.7 hours. This approach significantly enhances the accuracy-speed trade-off compared to CPU and GPU-based tools.
@@ 2.5 Time-Series Causal Discovery
The output of time series causal discovery algorithms are causal graphs [4]. These graphs could help researchers distinguish between causal and correlational relationships, support causal inference, and identify critical causal paths in complex systems. By visually representing causal relationships, it helps to understand and predict system behavior, guide experimental design and improve decision-making, leading to more accurate analysis and more reliable predictions.
@@ 2.6 Summary
This chapter introduces the basic concepts and methods for causal discovery in both non-temporal and time-series data, as well as techniques for analysing and accelerating these methods. First, it introduces time-series data to establish a basic understanding of dataset format. The chapter then explores causal discovery in non-temporal series data, starting with constraint-based approaches, which use conditions such as conditional independence tests to infer causality, and then function-based approaches, which model causal structure through functional relationships. After that, the chapter focuses on the causal discovery for time series data, using a similar approach but taking into account the time dependence and time lags in such data. Then, it discusses acceleration techniques, presenting strategies for optimizing constraint-based and function-based approaches to handle larger datasets and improve computational efficiency. Finally, the chapter summarises the differences between time-series causal discovery and non-temporal causal discovery, as well as the structure of three types of causal graphs, providing a foundation for subsequent discussions.
@@ 4.1 Speed Bottleneck of VarLiNGAM
VarLiNGAM [31] is a causal inference method for time series data with a process that combines the ideas of autoregressive modeling with LiNGAM. The workflow of VarLiNGAM is as follows:
@@ 4. Estimate the causal effect matrix $B_{\tau}$ by the following equation:
By using a performance analysis tool, it is possible to roughly analyze the time spent on each step throughout the model run. Among all these procedures, estimating the VAR coefficients and calculating the residual vectors only takes less than 1% of the total execution time, whereas the most time-consuming procedure is to perform a LiNGAM analysis to find the $B_{0}$ matrix, which takes over half of the total execution time when the dataset is large.
@@ 4.1.1 Time Complexity Analysis of LiNGAM
By using the performance analysis tool, it is possible to roughly analyze the time spent on each step throughout DirectLiNGAM model run. The main computational bottleneck of the DirectLiNGAM algorithm is _search_causal_order, which takes over 96% of the overall execution time.
@@ 4.2.1 Time Complexity After Precomputation
The idea of using pre-computed entropies in the _search_causal_order process is motivated by a common optimization strategy in algorithm design called avoiding redundant calculations. This approach is often referred to as memoization or precomputation, and it is particularly useful in scenarios where the same computationally expensive operations are repeatedly performed.
@@ 4.3 Summary
This chapter focuses on optimizing the VarLiNGAM algorithm by addressing its main speed bottlenecks and proposing acceleration strategies. First, it identifies that the most time-consuming process in VarLiNGAM is the LiNGAM analysis, specifically the step that searches for causal order. This step can take up a significant portion of the total execution time, especially for large datasets. To overcome this, the chapter introduces a precomputation strategy, which reduces redundant calculations by precomputing entropies and residual entropies before the causal order search. This significantly reduces the overall time complexity from $O(m^{3}\cdot n)$ to $O(m^{3}+m^{2}\cdot n)$, making the algorithm more efficient. Additionally, it also highlights the trade-off between time and space complexity, as precomputation increases memory usage. Nevertheless, these optimizations make VarLiNGAM still well-suited for large-scale datasets.
@@ 5.2.2 Validation: Fixed sample points
The graph 5.1(c) shows a sharp increase with larger feature sizes, similar to the trend seen in 5.1(a). This reinforces the observation that as feature size increases, the overall processing time becomes increasingly prohibitive without optimization techniques like precomputation.
@@ 5.2.3 Validation: Fixed feature size
As seen in 5.2(e) and 5.2(f), the performance of the causal graph improves slightly as the number of sample points increases and stabilises after about 500 sample points have been considered. Precomputation appears to have a stabilising effect on the performance metrics. This indicates that the algorithm may be more robust after optimization, maintaining its effectiveness across a broader range of data sizes. The consistency of the F1-score, Precision, Recall, and Accuracy after precomputation suggests that the capability of the algorithm is preserved despite the optimization.
@@ 5.3.1 Data from Simulation: fMRI
The minor reduction in the F1-score suggests that the precomputation does not significantly affect the performance of the model. This small difference is within the standard deviation range, indicating that the predictive performance remains stable even after optimization.
@@ 5.3.2 Real Dataset: IT Monitoring Datasets
For Web1, the F1-score decreases slightly after precomputation, indicating a minimal reduction in predictive performance. The decrease is very small, suggesting that the optimization did not significantly affect the model’s ability to classify or predict. The total execution time decreases from 2.588 seconds to 2.411 seconds after precomputation, despite the additional precomputation step. The significant reduction in the search causal order time (from 0.374 to 0.011 seconds) contributes to the overall efficiency improvement. The decrease in execution time is around 6.8%.
@@
The results for these datasets show that precomputation can significantly optimise time complexity without compromising, and in some cases slightly improving the performance. This makes precomputation a valuable optimization technique, especially for scenarios where large datasets are involved and quick execution is crucial.
@@ 6.1 Conclusion
This study addresses the challenges associated with causal discovery in time series data, particularly in the context of high-dimensional datasets. By optimizing the VarLiNGAM model, we have significantly reduced its computational complexity, making it feasible to handle large-scale datasets that were previously unmanageable.
Our work began with a comprehensive review of causal inference methods for both non-temporal and time series data, where we analyzed the performance of constraint-based and function-based approaches on complex datasets. We chose to focus on the VarLiNGAM model and designed a specialized dataset generator to generate simulated large-scale data. After that, we used some practical tools to identify the key bottlenecks in its computational processes. Through analysis of the whole process, we proposed acceleration methods tailored to these challenges by using precomputation on the entropies of each feature and the residuals of each pair. Specifically, we investigated the use of precomputation to streamline the execution of time-intensive functions, including _search_causal_order and _diff_mutual_info. The time complexity analysis shows that precomputation can significantly reduce the time complexity of the algorithm, a conclusion that is strongly supported by experimental results on a variety of datasets including simulated datasets, real fMRI datasets, IT monitoring datasets, and the SP500 financial dataset.